OpenStudy (anonymous):

Bill's father is building a ramp to reach the front porch of their house. The ramp will rise from the flat ground to the top of the porch. The porch is 2 feet tall. The ramp will rise at a 10 degree angle from the ground. How long does the ramp need to be? What is the sine of 10 degrees?

OpenStudy (anonymous):

 Remember that \[\sin\theta = \frac{opposite}{hypotenuse} ~\text{ so }~hypotenuse = \frac{opposite}{\sin\theta}\]